EXCITED STATES OF CH_{5}^{+} BY DIFFUSION MONTE CARLO

\underline{CHARLOTTE E. HINKLE}, ANNE B. McCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Understanding the spectrum of CH_{5}^{+} is a problem of long standing interest. Of particular interest is gaining a better understanding of the two high resolution spectra; one observed by White et. al ^{a}, and one that is partially assigned observed by the Nesbitt group. ^{b} Using excited state Diffusion Monte Carlo techniques in combination with symmetry adapted linear combinations of the internal coordinates of CH_{5}^{+}, obtained from group theoretical analysis of the G_{240} CNPI group, we investigate the nature of the vibrationally, and rotationally excited states of CH_{5}^{+}. By probing CH_{5}^{+} using these methods, we find that we are able to gain insights into the effects of the internal rotor motions of vibrationally excited states of CH_{5}^{+}.

aE. T. White, J. Tang, T. Oka, Science 284, 135 (1999).

bX. Huang., A. B. McCoy, J. M. Bowman, L. M. Johnson, C. Savage, F. Dong, D. J. Nesbitt, Science 311, 60 (2006).